An investor pays a NAV per share of €7.74 and purchases 100 shares in the Zebra Fund. After several months the NAV per share has dropped to €6.80 per share, however the fund still earned €0.75 per share in dividend income. What was the investors return on this investment?
Answer #1

Solution :

The return on investment is calculated using the following formula :

Return on Investment = [ ( Closing Price – Purchase Price + Dividend Income ) / Purchase Price ]

As per the Information given in the question we have

Closing Price = Closing NAV per share * No. of shares

= € 6.80 * 100 shares = € 680

Purchase Price = Purchase NAV per share * No. of shares

= € 7.74 * 100 shares = € 774

Dividend Income = $ 0.75 per share

Thus Total Dividend Income = Dividend Income per share* No. of shares

= € 0.75 * 100 =    € 75

Applying the above values in the formula we have

= ( 680 - 774 + 75 ) / 680

= - 19 / 680 = - 0.0279

= - 2.79 % ( when rounded off to two decimal places )

Thus Return on Investment = - 2.79 %